OpenStudy (anonymous):

three of four numbers have a sum of 22.if the average of the four number is 8,what is the number?

OpenStudy (anonymous):

\[\frac{1}{4} (a+b+c+d)=8 \]\[\frac{d+22}{4}=8 \]\[d=10 \]